Last year a manufacturer had a sales total for january that was 50 percent greater than the average of the monthly sales totals for February through December. The sales totals for january was what fraction of the sales total for the year?
Answer by stanbon(75887)

Last year a manufacturer had a sales total for january that was 50 percent greater than the average of the monthly sales totals for February through December. The sales totals for january was what fraction of the sales total for the year?
-------------
let "x" dollars be the average of monthly sales totals for Feb thru Dec.
Note: that is the average for 11 months.
Then the total of sales for the 11 months is 11x dollars
Then the sales for January was 1.5x.
So the Jan sales was 1.5x/(11x+1.5x)= 1.5x/12.5x = 1.5/12.5 = 3/25 of the
sales total for the year.
===============
cheers,
Stan H.